OpenStudy (anonymous):

how do i graph this inequality? lxl is less than or equal to 0. is it a ray, line segment,point, or line.

OpenStudy (mathmate):

Are you able to find any other number than zero that makes |x|\( le \) 0 ? If not then the solution set is a point.

OpenStudy (anonymous):

um...sorry im kinda confused

OpenStudy (mathmate):

x=0 satisfies the condition |x| <= 0?

OpenStudy (anonymous):

...\[\left| x \right| \le 0 \] and is it a point?? sorry its the first dsay of geomantry

OpenStudy (mathmate):

Exactly, since there is only one value of x that satisfies the requirement.

OpenStudy (anonymous):

Ohhh! i think i understand this now but how do i graph this on a number line?

OpenStudy (mathmate):



OpenStudy (mathmate):

Thank you! You're welcome! :)

OpenStudy (anonymous):

thank you so much!! you must take stats or something if you're a student :)

OpenStudy (mathmate):

I wish I knew more about stats! I guess one day I will.

OpenStudy (anonymous):

yes then i will understand geometry

OpenStudy (mathmate):

Yep, we'll both work on it!

OpenStudy (anonymous):

oh and 1 more question x+6y=40; A(-10,5) its just a subsitution but can you take a look at this please

OpenStudy (mathmate):

What are you supposed to do with the line and the point A?

OpenStudy (mathmate):

A is not on the line. Is it supposed to be?

OpenStudy (anonymous):

umm i the problem says : you are givin an equation of a line and a point. Use substitution to determine wheater or not the point is on the line

OpenStudy (mathmate):

Oops, I gave the answer away! This is how you do it: A(-10,5) means x=-10, y=5. Substitute these values on the left hand side to get (-10) + 6(5) = -10 + 30 = 20 \( \ne \) 40 therefore A is not on the line.

OpenStudy (anonymous):

ooops i accidently switched up the x and y intercepts....gotta do all the 6 problems all over again..

OpenStudy (mathmate):

Yep, (-10,5) means (x,y). They are kind of in alphabetical order. The good new: you'll remember it for a long time! :)

OpenStudy (anonymous):

yeah and i now just started my math semester i was so cluless with the practice problems we did

OpenStudy (mathmate):

You'll get back into it soon enough, as long as you keep practising and spend the time to work out problems as you do!

OpenStudy (anonymous):

yeah wish i found this site last semester then i might've passed Principles of Engineering with an A we did algebra geomenty trig like soh cah toa